OpenStudy (anonymous):

Find the coordinates of the points of intersection of the parabola y=x squared +3x-4 and the line y=2x+2. Find the coordinates of the points of intersection of the parabola y=x squared +3x-4 and the line y=2x+2. @Algebra

OpenStudy (asnaseer):

at the points where the two intersect, the x and y coordinates on both the parabola and the line must be equal. so, for the parabola we know:\[y=x^2+3x-4\]and for the line we know:\[y=2x+2\]we can therefore equate the right-hand-sides of both of these to find the x coordinate of the point of intersection. this gives us:\[x^2+3x-4=2x+2\]which can be re-arranged to:\[x^2+x-6=0\]this can fe factorised as follows:\[(x+3)(x-2)=0\]which gives us the x coordinates of the points of intersection as x=-3 and x=3. all you need to do now is to plug each of these x coordinates into either the equation for the parabola or the equation for the line to find what the y coordinates of the points of intersection are. the simplest is to use the equation of the line to find the y coordinates of the points of intersection.